### v1

# Collapse of the superconducting order parameter in Ising superconductors with Rashba spin-orbit coupling

Abstract
Ising superconductors have attracted quite some attention recently, due to their resilience against magnetic fields way beyond the Pauli-paramagnetic limit. Their protection against external magnetic field relies on strong Ising spin-orbit coupling, which originates from in-plane inversion symmetry breaking. Due to the heavy atom nature of Ising SCs, a smaller but sizable Rashba SOC could be present through gating or interfacial effects. Here, we consider the effect of Rashba SOC in a two valley model of Ising superconductors with an attractive $s$-wave interaction. We show that Rashba SOC gives a critical magnetic field, above which the superconducting order parameter collapses at low temperatures. This effect, however, disappears at high temperatures. Our findings demonstrate that the low- and high temperature physics of Ising SCs is quantitatively and qualitatively different in our two-valley model, and may lead to new ways to determine the strength of the Rashba SOC in Ising SCs.
Introduction.— Ising superconductors (SC) are doped two-dimensional transition metal dichalcogenides (TMDs), which are characterized by strong resilience against in-plane magnetic fields beyond the Pauli-paramagnetic limit. As a platform, TMDs are under active interest due to their direct band gap atomic scale thickness and strong spin-orbit coupling (SOC) [manzeli_2d_2017]. Like graphene, TMDs have a honeycomb lattice and hence two valleys. Unlike graphene, however, TMDs exhibit strong SOC because of their heavy atoms and in-plane inversion symmetry breaking. These features make them particularly interesting for potential applications in optoelectronics, valleytronics, and spintronics [xiao_coupled_2012, mak_control_2012]. When sufficiently doped, TMDs can become superconducting (SC) [kuzmanovic_tunneling_2022, de_la_barrera_tuning_2018, hamill_two-fold_2021, simon_transition-metal-dichalcogenide_2024, ilic_enhancement_2017, mockli_ising_2020, ramires_tailoring_2018, mockli_ising_2020, mockli_magnetic-field_2019, costanzo_tunnelling_2018]. Due to their large SOC, Cooper pairs are Ising coupled, making them resistant to large external magnetic fields [kuzmanovic_tunneling_2022, de_la_barrera_tuning_2018, hamill_two-fold_2021, simon_transition-metal-dichalcogenide_2024, ilic_enhancement_2017, mockli_ising_2020, ramires_tailoring_2018, mockli_ising_2020, mockli_magnetic-field_2019, costanzo_tunnelling_2018].
Besides the Ising SOC, one expects a Rashba SOC to originate from gating or interfacial effects in experiment—due to the heavy atoms of the Ising SC. The effect of Rashba SOC can lead to new physics. For instance, previously, it has been shown that the presence of Rashba SOC can lead to topological superconducting phases [yuan_possible_2014, meidan_josephson_junction_2024]. Experimentally, the strength of Rashba SOC has been claimed to be of the order of the SC order parameter [lu_evidence_2015, lu_full_2018], making Rashba SOC possibly important for understanding the phase diagram of Ising SCs.
In this Letter, we show that Rashba SOC sets a critical magnetic field—in the $s$-wave two valley model of Ising SCs—beyond which the superconducting order parameter (SC OP) collapses. This effect occurs at low temperatures and disappears close to the critical temperature $T_{c}$. While the effect of Rashba SOC is usually subtle, e.g., in the superconducting diode effect [yuan_topological_2021, ilic_theory_2022, khodas_diode_effect_Ising_2025]. This breakdown of the superconducting gap is a rather strong effect. Our results show that low and high temperature physic of SCs can behave both quantitatively and qualitatively different. Furthermore, our results may lead to new ways to determine the strength of the Rashba SOC in Ising SCs, which is currently a challenging task.
The model.— The low energy physics of doped transition metal dichalcogenides is described by two valleys around the $\mathrm{K}$ and $\mathrm{K}^{\prime}=-\mathrm{K}$ points and a gamma pocket $\Gamma$. In the proceeding, we focus on the physics of the $\pm\mathrm{K}$ points and ignore the physics of the Gamma pocket. The general Hamiltonian near the $\pm\mathrm{K}$ [xiao_coupled_2012, kormanyos_k_2015, mockli_ising_2020, ilic_enhancement_2017] points in the $(a_{\mathbf{p},\uparrow},a_{\mathbf{p},\downarrow})$ basis reads
$\hat{H}_{0}(\mathbf{p}+\eta\mathbf{K})=\left(\frac{\mathbf{p}^{2}}{2m^{*}}-\mu\right)\sigma_{0}+\eta\beta\sigma_{z}+\alpha(\hat{z}\times\mathbf{p})\cdot\boldsymbol{\sigma}+\mathbf{H}\cdot\boldsymbol{\sigma}$
with $\eta=\pm 1$ the valley number, $\boldsymbol{\sigma}$ the matrix of Pauli matrices and $\mathbf{H}$ the external magnetic field. The SOC coupling consists of an Ising SOC $\eta\beta\hat{z}$ and a Rashba SOC $\alpha(\hat{z}\times\mathbf{p})$, see Fig. 1. We consider the interactions to be an attractive four point interaction towards $s$-wave superconductivity. Consequently, the mean-field Bogoliubov-de-Gennes (BdG) Hamiltonian in the basis $\big(a_{\mathbf{k},\uparrow},a_{\mathbf{k},\downarrow},a_{\mathbf{-k},\uparrow}^{\dagger},a_{\mathbf{-k},\downarrow}^{\dagger}\big)$ becomes
$\hat{H}_{\mathrm{BdG}}=\begin{pmatrix}H_{0}(\mathbf{k})&\mathrm{i}\sigma_{y}\Delta\\ -\mathrm{i}\sigma_{y}\Delta^{*}&-H_{0}^{*}(\mathbf{-k})\end{pmatrix}$, (1)
with $\mathbf{k}=\mathbf{p}+\eta\mathbf{K}$ and denotes complex conjugation. Moreover, $\Delta$ is the SC singlet order parameter (OP). From the BdG Hamiltonian, the Gorkov Green’s function is defined as $\check{{\mathcal{G}}}_{\mathrm{i}\omega_{n}}^{-1}(\mathbf{k})=\mathrm{i}\omega_{n}-\hat{H}_{\mathrm{BdG}}.$
To simplify calculations, we diagonalize the normal state in spin space using the unitary transformation $\hat{\mathcal{U}}_{\mathbf{k}}=\hat{U}_{\mathbf{k}}\oplus\hat{U}^{*}_{\mathbf{-k}},$ with $\hat{U}_{\mathbf{k}}=e^{-{\mathrm{i}}\phi_{\mathbf{k}}\sigma_{z}/2}e^{-\mathrm{i}\theta_{\mathbf{k}}\sigma_{y}/2}.$ The angles used above are defined via $\sin(\theta_{\mathbf{k}})\cos(\phi_{\mathbf{k}}),\sin(\theta_{\mathbf{k}})\sin(\phi_{\mathbf{k}}),\cos(\theta_{\mathbf{k}})]={(\eta\beta\hat{z}+\alpha(\hat{z}\times\mathbf{p})+\mathbf{H})}/{\|\eta\beta\hat{z}+\alpha(\hat{z}\times\mathbf{p})+\mathbf{H}\|}$. The SC OP transforms equivalently and reads $\Delta\mathrm{i}\sigma_{y}\rightarrow\Delta\hat{U}_{\mathbf{k}}^{\dagger}\mathrm{i}\sigma_{y}\hat{U}^{*}_{\mathbf{-k}}=\Delta(\bar{\beta}\sigma_{0}+\bar{H}\mathrm{i}\sigma_{y})+\mathcal{O}(\alpha^{2}),$ with $\bar{\beta}={\beta}/{\sqrt{H^{2}+\beta^{2}}}$ and $\bar{H}=H/\sqrt{H^{2}+\beta^{2}}$. By grouping together spin-up and spin-down sectors in the BdG Hamiltonian (1), the inverse Green’s function splits into a BCS-like sector with a rescaled SC OP $\Delta\rightarrow\bar{\beta}\Delta$ and an off-diagonal sector governed by the external magnetic field strength $\bar{H}$
$\check{\mathcal{G}}_{\mathrm{i}\omega_{n}}^{-1}(\mathbf{k})=\mathrm{i}\omega_{n}\mathbb{I}$ $-\bigoplus_{\nu=\pm}\begin{pmatrix}\epsilon_{\nu}(\mathbf{k})&\bar{\beta}\Delta\\ \bar{\beta}^{*}\Delta^{*}&-\epsilon_{\nu}(\mathbf{-k})\end{pmatrix}-\mathrm{i}\sigma_{y}\otimes\delta\check{H},$
where we introduced $\delta\check{H}=\begin{pmatrix}0&\bar{H}\Delta\\ -\bar{H}^{*}\Delta^{*}&0\end{pmatrix}$ for notational convenience.
Figure 1: Graphical depiction of the in-plane component of the spin-splitting at the Fermi surface due to Rashba SOC and an external magnetic field in the $x$-direction. The out of plane spin splitting is given by the Ising SOC.
The off-diagonal term $\mathrm{i}\sigma_{y}\otimes\delta\check{H}$ in the inverse Green’s function above, can be treated perturbatively since $|\Delta|/\sqrt{H^{2}+\beta^{2}}$ is found to be at most of the order $\mathcal{O}(10^{-1})$ in Ising SCs [de_la_barrera_tuning_2018, falson_type-ii_2020, hamill_two-fold_2021, xi_ising_2016, kuzmanovic_tunneling_2022, lu_full_2018, lu_evidence_2015]. Hence, to good accuracy we may expand the Gorkov Green’s function up to first order. The lowest order Gorkov Green’s function has a BCS form and reads
$\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},\nu}\simeq$ $\frac{\begin{pmatrix}\mathrm{i}\omega_{\nu}+\epsilon_{\nu}(\mathbf{-k})&\bar{\beta}\Delta\\ \bar{\beta}^{*}\Delta^{*}&\mathrm{i}\omega_{n}-\epsilon_{\nu}(\mathbf{k})\\ \end{pmatrix}}{(\mathrm{i}\omega_{n}-\epsilon_{\nu}(\mathbf{k}))(\mathrm{i}\omega_{n}+\epsilon_{\nu}(\mathbf{-k}))-\bar{\beta}^{2}|\Delta|^{2}},$
with $\epsilon_{\nu}(\pm\mathbf{k})=\mathbf{p}^{2}/2m^{*}+\nu\sqrt{H^{2}+\beta^{2}}\pm\alpha\nu(\mathbf{\bar{H}}\times\hat{z})\cdot\mathbf{p}$. We obtain the contribution from the external magnetic field by expanding the Gorkov Green’s function up to first order in $\mathrm{i}\sigma_{y}\otimes\delta\check{H}$,
$\check{\mathcal{G}}_{\mathrm{i}\omega_{n}}\simeq$ $\begin{pmatrix}\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},+}&-\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},+}\delta H\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},-}\\ -\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},-}\delta H^{\dagger}\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},+}&\check{\mathcal{G}}^{0}_{\mathrm{i}\omega_{n},-}\end{pmatrix}$, (2)
which gives the Green’s function up to the desired accuracy. Since our goal is to determine the self-consistency equation, we only need the anomalous Green’s function. Within the quasi-classical approximation, the diagonal part of the anomalous Green’s function in Eq. 2 reads
$\frac{f_{\mathrm{i}\omega_{n},\nu,\nu}^{0}(\eta,\phi)}{\bar{\beta}\Delta}=\frac{2\pi\mathrm{i}\mathcal{N}}{\mathrm{i}\sqrt{\left(\omega_{n}+\mathrm{i}\big[\alpha\nu(\mathbf{\bar{H}}\times\hat{z})\cdot\mathbf{p}_{F}\big]\right)^{2}+\bar{\beta}^{2}|\Delta|^{2}}},$
while the off-diagonal, magnetic field dependent, contributions give
$\frac{f_{\mathrm{i}\omega_{n},+,-}^{0}(\eta,\phi)}{\bar{H}\Delta}=$ $\sum_{\nu}\frac{2\pi\mathrm{i}\mathcal{N}\mathrm{sign}(\omega_{n})/2}{\mathrm{i}\omega_{n}-\sqrt{H^{2}+\beta^{2}}},$
$\frac{f_{\mathrm{i}\omega_{n},-,+}^{0}(\eta,\phi)}{-\bar{H}\Delta}=$ $\sum_{\nu}\frac{2\pi\mathrm{i}\mathcal{N}\mathrm{sign}(\omega_{n})/2}{\mathrm{i}\omega_{n}+\sqrt{H^{2}+\beta^{2}}}.$
We defined $\mathcal{N}$ to be the electron density of states in the normal state.
From here, we are in the position to write down the self-consistency equation. In the spin-diagonal basis, this reads
$\frac{\Delta}{\lambda}=\frac{k_{B}T}{2}\sum_{n}\int\frac{\mathrm{d}\phi}{2\pi}\mathrm{Tr}\left\{f_{\mathrm{i}\omega_{n}}\left(\mathcal{U}_{\mathbf{-k}}^{T}\mathrm{i}\sigma_{y}\mathcal{U}_{\mathbf{k}}\right)\right\}.$
Here, $\lambda$ denotes the magnitude of the attractive four point interaction, which relates directly to the SC OP at zero temperature $1/\mathcal{N}\lambda=\log(2\omega_{D}/|\Delta_{0}|)$, with $\omega_{D}$ the Debye frequency. Accordingly, the self-consistency equation for the SC OP becomes
$\frac{1}{\mathcal{N}\lambda}=\pi k_{B}T\sum_{n,\nu}\int\frac{\mathrm{d}\phi}{2\pi}\;\Bigg[\frac{\bar{H}^{2}\mathrm{sign}(\omega_{n})}{\omega_{n}+\mathrm{i}\nu\sqrt{H^{2}+\beta^{2}}}$ (3)
$+\frac{\bar{\beta}^{2}}{\sqrt{\left(\omega_{n}+\mathrm{i}\big[\alpha\nu(\mathbf{\bar{H}}\times\hat{z})\cdot\mathbf{p}_{F}\big]\right)^{2}+\bar{\beta}^{2}|\Delta|^{2}}}\Bigg]$ .
In the following, we discuss the effect of Rashba SOC on the magnetic field dependence of the SC OP at low temperatures. In principle, the effect of Rashba on Ising SCs is two-fold. One effect is that Rashba SOC coupling alters the thermodynamic ground state to a helical state $\Delta\rightarrow\Delta_{0}e^{\mathrm{i}qx}$ [kaur_helical_2005, yuan_topological_2021]. This effect is however negligible with respect to the magnitude of the SC OP since $v_{F}q$ is at most of the order $\mathcal{O}(\alpha p_{F}/\beta)$ and the correction of the helical phase to $|\Delta|/\Delta_{0}$ is at most of the order $\mathcal{O}(\alpha^{2}p_{F}^{2}/\beta^{2})<\mathcal{O}(10^{-2})$. However, the main effect of Rashba SOC to the magnitude of the SC OP is that the SC OP collapses beyond a critical magnetic field set by Rashba SOC, as we will show next.
Collapse of the SC OP due to Rashba SOC at low temperatures.—
Figure 2: The self-consistent superconducting order parameter $\bar{\beta}|\Delta|$ as a function of the external magnetic field $H$ for different values of Rashba SOC $\alpha p_{F}$ at the temperature $T/T_{c}=10^{-2}$ and Ising SOC strength $\beta=7\Delta_{0}$. The effect of Rashba SOC is clear, as the gap reaches $\alpha p_{F}H=\beta|\Delta|$ it collapses.
In the following, we focus on low temperatures $k_{B}T\ll\alpha p_{F}\lesssim\Delta_{0}\ll\sqrt{H^{2}+\beta^{2}}$. Translated to the critical temperature, this means $T/T_{c}\lesssim 10^{-2}$. Since $T_{c}\sim\mathcal{O}(10\mathrm{K})$ in Ising SCs, this corresponds to temperatures of the order $T\lesssim\mathcal{O}(100\mathrm{mK})$ [kuzmanovic_tunneling_2022, hamill_two-fold_2021, li_recent_2021, xi_ising_2016]. Such temperatures are typically accessible in experiment [kuzmanovic_tunneling_2022]. Furthermore, we consider the energy scale of Rashba SOC to be at most of the order of the energy scale of the SC OP at zero temperature. As detailed in the supplementary material, we perform the summation over Matsubara frequencies in the first term of Eq. 3, which is known to give the digamma function [kopnin_theory_2009] and approximate the second term for low temperatures. As a result, the self-consistency equation (3) for low temperatures reads
$\log\left(\frac{\bar{\beta}|\Delta|}{\Delta_{0}}\right)+\frac{H^{2}}{\beta^{2}}\left[\log\left({2\sqrt{H^{2}+\beta^{2}}}/{\Delta_{0}}\right)\right]$ (4)
$=-\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right).$
The left-hand side of Eq. 4 describes the magnetic field dependence of the Ising SC in the absence of Rashba SOC. The right-hand side can be separated into two regions. If $\alpha p_{F}H\lesssim\beta|\Delta|$, it is negligibly small and the SC OP does not change quantitatively due to the presence of Rashba SOC. However, when $\alpha p_{F}H\gtrsim\beta|\Delta|$, it dominates the self-consistency equation and causes the SC OP to collapse.
The magnetic field dependence of the SC OP in Eq. 4 can be solved for low temperatures and gives
$\frac{\bar{\beta}|\Delta|}{\Delta_{0}}\simeq\left(\frac{\Delta_{0}}{2\sqrt{H^{2}+\beta^{2}}}\right)^{{H^{2}}/{\beta^{2}}}\Theta\left(\beta|\Delta|-\alpha p_{F}H\right),$ (5)
with $\Theta$ the Heaviside step function being the low temperature limit of the Fermi-Dirac distribution $n_{F}\left[({\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|})/{k_{B}T}\right]$. This is the main result of this Letter. Namely, the superconducting gap obtains a cutoff due to Rashba SOC given by $\alpha p_{F}H\sim\beta\Delta$, as can be seen in Fig. 2. This collapse happens precisely when the energy of Rashba SOC times magnetic field strength equals that of the superconducting gap times Ising SOC. As we show later, this corresponds to the point at which the density of states of quasi-particles at zero energy becomes non-vanishing. Due to the non-linear nature of this collapse, it disappears when going to the high temperature Ginzburg-Landau limit $|\Delta|\ll k_{B}T$, see supplementary material. This shows how the high- and low temperature physics of Ising SC can be quantitatively and qualitatively different.
For completeness, we determine the critical magnetic field due to Rashba SOC. We first note that Eq. 5 can be approximated further by $|\Delta|/\Delta_{0}\sim\left(\Delta_{0}/2\beta\right)^{{H^{2}}/{\beta^{2}}}$. The critical SC OP size is given by $\alpha p_{F}H=\beta\Delta$, which together with the previous equality gives $\alpha p_{F}H/\beta=\Delta_{0}\left(\Delta_{0}/2\beta\right)^{{H^{2}}/{\beta^{2}}}$. We find the critical magnetic field to be given by ${H_{\alpha}^{2}}/{\beta^{2}}={-W[-2(\Delta_{0}/\alpha p_{F})^{2}\log(\Delta_{0}/2\beta)]}/{2\log(\Delta_{0}/2\beta)},$ with $W$ the Lambert $W$ function. We show the dependence of the critical magnetic field as a function of Rashba SOC in Fig. 3. We find the effect of Rashba SOC to be significant in the superconducting phase diagram for $\alpha p_{F}/\Delta_{0}\gtrsim\mathcal{O}(10^{-1})$.
Figure 3: Dependence of the critical magnetic field on the strength of Rashba for $\beta=7\Delta_{0}$. We notice that Rashba SOC is still significant if it is one order smaller than the superconducting gap $\alpha p_{F}/\Delta_{0}\gtrsim\mathcal{O}(10^{-1})$.
In Refs. [lu_evidence_2015, lu_full_2018] values of $\alpha p_{F}\sim\Delta_{0}$ in $\mathrm{MoS_{2}}$ and $\mathrm{WS_{2}}$ were reported. This should be sufficient for the collapse of the SC OP to happen at low temperatures.
Typically, the SC OP is measured via the density of states using tunneling techniques. As we will show, the effective SC gap in the density of states differs from $|\Delta|$ by a factor of $\sqrt{1+H^{2}/\beta^{2}}$. Up to the precision of the Gorkov Green’s function in Eq. 2, the density of states is given by
$\mathcal{N}_{s}(\epsilon)=\mathcal{N}\mathrm{Re}\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{2\pi}\,\frac{\epsilon+\alpha p_{F}\bar{H}\cos(\phi)}{\sqrt{(\epsilon+\alpha p_{F}\bar{H}\cos(\phi))^{2}-\bar{\beta}^{2}|\Delta|^{2}}}.$
Hence, the effective gap that is measured via the density of states is $\bar{\beta}|\Delta|$ rather than $|\Delta|$ and Rashba SOC makes the density of states cone-shaped for energies $\epsilon$ satisfying $|\epsilon-\bar{\beta}|\Delta||<\alpha p_{F}H$, see Fig. 4. Moreover, inclusion of the second order correction in Gorkov Green’s function (2) gives what is called a mirage gap at high energies $\epsilon\sim\sqrt{\beta^{2}+H^{2}}$ in the density of states [tang_magnetic_2021].
Figure 4: Density of states of the Ising SC for different values of $\alpha p_{F}H/\beta|\Delta|$. This describes BCS density of states for $\alpha p_{H}=0$ with a gap given by $\bar{\beta}|\Delta|$ and becomes cone-shaped within $|\epsilon-\bar{\beta}|\Delta||<\alpha p_{F}H$ for increasing values of $\alpha p_{F}H/\beta|\Delta|$. For $\alpha p_{F}H=\beta|\Delta|$ the SC OP collapses for low temperatures, which gives a constant density of states. For high temperatures, however, the SC OP doesn’t collapse beyond the critical magnetic field, rendering the density of states gapless and cone-shaped for higher magnetic fields.
Discussion.— In this Letter, we showed that the presence of Rashba SOC in Ising superconductors can strongly affect the magnetic field dependence of the superconducting order parameter at low temperatures. Our main finding is that the SC OP collapses beyond a critical magnetic field satisfying $\alpha p_{F}H=\beta|\Delta|$. This collapse disappears at high temperatures—close to $T_{c}$—and demonstrates how the low- and high temperature limits of Ising SCs can be quantitatively and qualitatively different. We find Rashba SOC to be significant to the superconducting phase diagram if $\alpha p_{F}/\Delta_{0}\gtrsim\mathcal{O}(10^{-1})$. Due to the heavy atom nature of TMDs, such a Rashba SOC could be induced from gating or interfacial effects in experiments. Experimentally, Rashba SOCs of magnitude of the SC OP have already been measured in thin-film Ising SCs [lu_evidence_2015, lu_full_2018]. Our findings may lead to new ways to determine the size of Rashba SOC in Ising SCs, which is currently a challenging task.
The collapse of the SC OP at low temperatures should be experimentally verifiable using tunneling spectroscopy [kuzmanovic_tunneling_2022, dvir_spectroscopy_2018, costanzo_tunnelling_2018]. While different tunneling devices have different sensitivities towards either the $\Gamma$ pocket or the $\pm\mathrm{K}$ valleys, the predicted effect should be within experimental reach using tunneling devices that are mostly sensitive to the density of states in the $\pm\mathrm{K}$ valleys [kuzmanovic_tunneling_2022].
In this work, we considered the $s$-wave two valley model of Ising SCs and disregarded the influence of the $\Gamma$ pocket. This is justified for measurement devices sensitive to the valleys and for high magnetic fields—since Ising SOC is larger in the valleys. Furthermore, we considered a clean Ising SC and assumed tunneling between the $\Gamma$ pocket and the $\pm\mathrm{K}$ valleys to be negligible.
In future work, one could extend this model to include the $\Gamma$ pocket. Additionally, inclusion of impurities and scattering between the $\pm\mathrm{K}$ valleys and $\Gamma$ pocket would be of interest.
Acknowledgements
JSH would like to thank P. M. Gunnink for input on the manuscript. We acknowledge financial support by Deutsche Forschungsgemeinschaft (DFG, German Re-search Foundation)–Project-ID 443404566 - SPP 2244.
I Self-consistency equation in the low temperature limit
The self consistency equation from the original basis to the rotated basis is given by
$\frac{\Delta}{\lambda}=$ $\frac{k_{B}T}{2}\sum_{n}\int\frac{\mathrm{d}\phi}{2\pi}\mathrm{Tr}\left\{f_{\mathrm{i}\omega_{n}}\mathrm{i}\sigma_{y}\right\}\rightarrow\frac{k_{B}T}{2}\sum_{n}\int\frac{\mathrm{d}\phi}{2\pi}\mathrm{Tr}\left\{f_{\mathrm{i}\omega_{n}}\left(\mathcal{U}_{\mathbf{-k}}^{T}\mathrm{i}\sigma_{y}\mathcal{U}_{\mathbf{k}}\right)\right\}.$ (6)
Filling in the anomalous Green’s function, we find the self-consistency equation (3), which reads
$\frac{1}{\mathcal{N}\lambda}=\pi\mathrm{i}k_{B}T\sum_{n,\nu}$ $\int\frac{\mathrm{d}\phi}{2\pi}\left[\frac{\bar{\beta}^{2}}{\mathrm{i}\sqrt{\left(\omega_{n}+\mathrm{i}\big[\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F,\nu}\big]\right)^{2}+\bar{\beta}^{2}|\Delta|^{2}}}+\frac{\bar{H}^{2}\mathrm{sign}(\omega_{n})}{\mathrm{i}\omega_{n}-\nu\sqrt{H^{2}+\beta^{2}}}\right].$ (7)
We furthermore assume $k_{B}T\ll\alpha p_{F}\lesssim\Delta_{0}\ll\sqrt{H^{2}+\beta^{2}}$ and notice that the sum over Matsubara frequencies of the left term can be transformed into an integral in real space
$\frac{1}{\mathcal{N}\lambda}=$ (8)
$\sum_{\nu}\int\frac{\mathrm{d}\epsilon}{4}\int\frac{\mathrm{d}\phi}{2\pi}\;\frac{\bar{\beta}^{2}\tanh\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F,\nu}\right)}/{2k_{B}T}\right]}{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}+\pi\mathrm{i}k_{B}T\sum_{n,\nu}\Bigg[\frac{\bar{H}^{2}\mathrm{sign}(\omega_{n})}{\mathrm{i}\omega_{n}-\nu\sqrt{H^{2}+\beta^{2}}}\Bigg].$
In order to perform the integral in Eq. 8, we use the identity $\tanh(x)=1-2n_{F}(x)$. By noticing that $1/\mathcal{N}\lambda=\bar{\beta}^{2}/\mathcal{N}\lambda+\bar{H}^{2}/\mathcal{N}\lambda$, we may write the first line in Eq. 8 as
$\frac{1}{\mathcal{N}\lambda}+\sum_{\nu}\int\frac{\mathrm{d}\epsilon}{4}\frac{1}{\sqrt{\epsilon^{2}+\bar{\beta}|\Delta|^{2}}}-\int\frac{\mathrm{d}\epsilon}{2}\frac{\mathrm{d}\phi}{2\pi}\;\frac{n_{F}\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}\right)}/{k_{B}T}\right]}{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}.$ (9)
The first two terms in Eq. 9 give $\log(\Delta_{0}/\bar{\beta}|\Delta|)$. The integral over the Fermi-distribution on the right-hand side of Eq. 9 may be rewritten as a series expansion
$\int\frac{\mathrm{d}\phi}{2\pi}n_{F}\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}\right)}/{k_{B}T}\right]$ (10)
$=$ $\int\frac{\mathrm{d}\phi}{2\pi}\sum_{n=1}^{\infty}(-1)^{n-1}\exp\left[-n\frac{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}}{k_{B}T}\right]$
$=$ $\sum_{n=1}^{\infty}(-1)^{n-1}\exp\left[-n\frac{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}{k_{B}T}\right]I_{0}\left(n\frac{\alpha p_{F}\bar{H}}{k_{B}T}\right).$
For $\alpha p_{F}\gg k_{B}T$ the modified Bessel function becomes $I_{0}\left({n\alpha p_{F}\bar{H}}/{k_{B}T}\right)\simeq\sqrt{{k_{B}T}/{2\pi n\alpha p_{F}\bar{H}}}\exp\left({n\alpha p_{F}\bar{H}}/{k_{B}T}\right).$ Similarly, for $|\Delta|\gg k_{B}T$ we may approximate the exponential in Eq. 10 by ${\exp\left(-n{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}/{k_{B}T}\right)}/{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}\simeq{\exp\left(-n{(\bar{\beta}|\Delta|+\epsilon^{2}/2\bar{\beta}|\Delta|)/k_{B}T}\right)}/{\bar{\beta}|\Delta|}.$ Using these approximations, the integral over energy in Eq. 8 reads
$\int\frac{\mathrm{d}\epsilon}{2}\frac{\mathrm{d}\phi}{2\pi}\;\frac{n_{F}\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}\right)}/{k_{B}T}\right]}{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}},}$ (11)
$\simeq\sum_{n=1}^{\infty}\int\frac{\mathrm{d}\epsilon}{2}\;(-1)^{n-1}\frac{\exp\left[-\frac{n\epsilon^{2}}{2\bar{\beta}|\Delta|k_{B}T}\right]}{\bar{\beta}|\Delta|}\sqrt{\frac{k_{B}T}{2\pi n|\alpha|}}\exp\left[\frac{n(\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|)}{k_{B}T}\right]$
$=\sum_{n=1}^{\infty}\frac{(-1)^{n-1}}{2n}\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\exp\left[\frac{n(\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|)}{k_{B}T}\right]$
$=\frac{1}{2}\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right).$
Moreover, the sums over Matsubara frequencies on the right-hand hand side of Eq. 8 are known to give polygamma functions. To obtain the final expression for the self-consistency equation, we reinterpret the BCS energy cutoff $\omega_{D}$ in terms of $T_{c}$ and $\Delta_{0}$. To achieve this, we use that $1/\mathcal{N}\lambda=\log(2\omega_{D}/|\Delta_{0}|)=\sum_{n=0}^{\hbar\omega_{D}/2\pi k_{B}T_{c}}{(n+1/2)^{-1}}$ at zero field. Hence, the self-consistency equation at low temperatures $\Delta\gg k_{B}T$ simplifies to
$\log\left(\frac{|\Delta_{0}|}{\bar{\beta}|\Delta|}\right)=$ $\sum_{\nu}\frac{1}{2}\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right)+\frac{H^{2}}{\beta^{2}}\Bigg[\log\left(\frac{T}{T_{c}}\right)-\psi\left(\frac{1}{2}\right)+\log\left(\frac{\sqrt{H^{2}+\beta^{2}}}{2\pi k_{B}T}\right)\Bigg].$ (12)
[…]
$\log\left(\frac{|\Delta_{0}|}{\bar{\beta}|\Delta|}\right)=$ $\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right)+\frac{H^{2}}{\beta^{2}}\log\left(\frac{2\sqrt{H^{2}+\beta^{2}}}{\Delta_{0}}\right).$ (13)
II High temperature Ginzburg-Landau limit
To determine the high temperature limit, $(T_{c}-T)/T\ll 1$, we expand the self-consistency equation (7) up to second order in $\bar{\beta}|\Delta|$ and $\alpha p_{F}$ and ignore the contribution from the helical state $v_{F}q\rightarrow 0$. After taking the angular integral, the self-consistency equation (7) reads
$\frac{1}{\mathcal{N}\lambda}=\pi\mathrm{i}k_{B}T\sum_{n,\nu}\;\bar{\beta}^{2}\mathrm{sign}(\omega_{n})$ $\left[\frac{1}{\mathrm{i}{\omega_{n}}}-\frac{\alpha^{2}p_{F}^{2}\bar{H}^{2}+\bar{\beta}^{2}|\Delta|^{2}}{2\mathrm{i}\omega_{n}^{3}}\right]$ (14)
$+{\bar{H}^{2}\mathrm{sign}(\omega_{n})}$ $\left[\frac{1}{\mathrm{i}\omega_{n}-\nu\sqrt{H^{2}+\beta^{2}}}\right].$
The sums over Matsubara frequencies are known to give polygamma functions and the above self-consistency equation becomes
$\log\left(\frac{T}{T_{c}}\right)=\bar{H}^{2}$ $\left[\psi\left(\frac{1}{2}\right)-\mathrm{Re}\psi\left(\frac{1}{2}-\mathrm{i}\frac{\sqrt{H^{2}+\beta^{2}}}{2\pi k_{B}T}\right)\right]-\bar{\beta}^{2}\left[\frac{7\zeta(3)}{2(2\pi k_{B}T)^{2}}\left(\alpha^{2}p_{F}^{2}\bar{H}^{2}+\bar{\beta}^{2}|\Delta|^{2}\right)\right].$ (15)
Since the Ising SOC is typically much larger than the critical temperature $\sqrt{\beta^{2}+H^{2}}\gg k_{B}T$ the digamma function is well approximated by a logarithm $\mathrm{Re}\psi\left({1}/{2}-\mathrm{i}{\sqrt{H^{2}+\beta^{2}}/}{2\pi k_{B}T}\right)\rightarrow\log\left({\sqrt{H^{2}+\beta^{2}}}/{2\pi k_{B}T}\right).$ Thus, the above is well approximated by
$\log\left(\frac{T}{T_{c}}\right)=\bar{H}^{2}\left[\psi\left(\frac{1}{2}\right)-\log\left(\frac{\sqrt{H^{2}+\beta^{2}}}{2\pi k_{B}T}\right)\right]-\bar{\beta}^{2}$ $\left[\frac{7\zeta(3)}{2(2\pi k_{B}T)^{2}}\left(\alpha^{2}p_{F}^{2}\bar{H}^{2}+\bar{\beta}^{2}|\Delta|^{2}\right)\right].$ (16)
The above can be written more compactly as
$\frac{7\zeta(3)}{2(2\pi k_{B}T)^{2}}\left(\bar{\beta}^{2}|\Delta|^{2}+\alpha^{2}p_{F}^{2}\bar{H}^{2}\right)=\log\left(\frac{T_{c}}{T}\right)-\frac{H^{2}}{\beta^{2}}\log\left(\frac{2\sqrt{H^{2}+\beta^{2}}}{\Delta_{0}}\right).$ (17)
Hence, we see the SC OP does not collapse at $\beta|\Delta|=\alpha p_{F}H$ for temperatures close to $T_{c}$.